OpenStudy (anonymous):

Find the equation of the line that has a slope of 1/4 and passes through (8, -3)

OpenStudy (gorv):

use y-y1=slope(x-x1) and (x1,y1)=(8,-3) implies x1=8 and y1=-3

OpenStudy (anonymous):

im still lost so what would the answer be?

OpenStudy (anonymous):

??

OpenStudy (whpalmer4):

Use the formula for a line with known slope, \(m\), passing through known point \((x_0,y_0)\): \[y-y_0 = m(x-x_0)\] You know \(m = \frac{1}{4}\) and \((x_0,y_0)\) is \((8,-3)\) Plug in the values to get your line equation. Be careful about subtracting negative numbers...

OpenStudy (gorv):

y+3=(1/4)*(x-8)

OpenStudy (anonymous):

\[y+3=\frac{ 1 }{ 4}(x-8)\]

OpenStudy (gorv):

yes

OpenStudy (anonymous):

thanks

OpenStudy (whpalmer4):

Might want to rearrange it into slope-intercept form: \(y = mx + b\) but you do have a valid equation for that line, as you can see from the plot here:

OpenStudy (gorv):

say thanks to @whpalmer4 also

OpenStudy (whpalmer4):

I figured it was a thanks to the group :-)

OpenStudy (anonymous):

it was for both of you guys :)

OpenStudy (whpalmer4):

Happy to be of some help! But now it's time to hit the sack for me...